Skip to main contentSkip to solution

How many minimum number of times must a man toss a fair coin so that the probability of having at least one head is more than 90%?

Solution

✅ Correct Option: 2

The probability of getting at least one head can be calculated using the complement:

P(at least one head) = 1 - P(all tails)

For a fair coin, P(tail) = 12\frac{1}{2}


If the coin is tossed nn times:

P(all tails) = (12)n\left(\frac{1}{2}\right)^n

P(at least one head) = 1−(12)n1 - \left(\frac{1}{2}\right)^n


The probability must be more than 90%:

1−(12)n>0.901 - \left(\frac{1}{2}\right)^n > 0.90

(12)n<0.10\left(\frac{1}{2}\right)^n < 0.10


For n=3n = 3:

(12)3=18=0.125\left(\frac{1}{2}\right)^3 = \frac{1}{8} = 0.125

P(at least one head) = 1−0.125=0.875=87.5%1 - 0.125 = 0.875 = 87.5\%

This is less than 90%.


For n=4n = 4:

(12)4=116=0.0625\left(\frac{1}{2}\right)^4 = \frac{1}{16} = 0.0625

P(at least one head) = 1−0.0625=0.9375=93.75%1 - 0.0625 = 0.9375 = 93.75\%

This is more than 90%.


Therefore, the minimum number of tosses required is 4.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question